OpenStudy (anonymous):

find the mid point of the segment having the given endpoints (5, -10) and (-15, -2) Not looking to cheat, this is just a question from a review sheet. the test is tomorrow, and if you could provide the asnwer with the proccess of HOW to solve this so i don't fail, it'd be very much appreciated. thank you!

OpenStudy (anonymous):

Use the midpoint formula. \(\huge{(} \normalsize\dfrac{x_1 + x_2}{2}, \dfrac{y_1 + y_2}{2}\huge)\)

OpenStudy (anonymous):

Plug in the points: \(\huge{(} \normalsize\dfrac{5 - 15}{2}, \dfrac{-10 - 2}{2}\huge)\) Simplify, what do you get? @xHistoryMajorx

OpenStudy (anonymous):

@xHistoryMajorx Can you subtract 5 - 15?